Bay City Company’s fixed budget performance report for July follows. The $440,000 budgeted total expenses include $300,000 variable expenses and $140,000 fixed expenses. Actual expenses include $130,000 fixed expenses. Fixed Budget Actual Results Variances Sales (in units) 6,000 4,900 Sales (in dollars) $ 480,000 $ 431,200 $ 48,800 U Total expenses 440,000 406,000 34,000 F Income from operations $ 40,000 $ 25,200 $ 14,800 U Prepare a flexible budget performance report that shows any variances between budgeted results and actual results. List fixed and variable expenses separately.
Answer #1

Answer-

BAY CITY COMPANY
FLEXIBLE BUDGET PERFORMANCE REPORT
FOR THE MONTH ENDED
PARTICULARS FLEXIBLE BUDGET ACTUAL RESULTS VARIANCES REMARK
$ $ $
Sales ($480000/6000 units)*4900 units =392000 431200 39200 Favorable
Less- Variable Expenses ($300000/6000 units)*4900 units =245000 406000-130000 =276000 -31000 Unfavorable
Contribution margin 147000 155200 8200 Favorable
Less- Fixed Expenses 140000 130000 10000 Favorable
Income from operations 7000 25200 18200 Favorable